Can you solve the system of equation by substitution?
x-8y=-23
5y-2x=2

Answer by Maths68(1474)

x-8y=-23.....(1)
5y-2x=2......(2)
Take (1)
x-8y=-23
x=8y-23......(3)
Put the value of x in (2)
5y-2(8y-23)=2
5y-16y+46=2
5y-16y=2-46
-11y=-44
-11y/-11=-44/-11
y=4

Put the value of y in (2)
5y-2x=2......(2)
5(4)-2x=2
20-2x=2
-2x=2-20
-2x=-18
-2x/-2=-18/-2
x=9
Solution set (x,y)=(9,4)
Check
=====
Put the values of x and y in (1)
x-8y=-23.....(1)
9-8(4)=-23
9-32=-23
-23=-23
Put the values of x and y in (2)
5y-2x=2......(2)
5(4)-2(9)=2
20-18=2
2=2